Two equal positive point charges 𝑞1 and 𝑞2 = 4.0 µC interact with point charge Q= 5.0 µC. Find the magnitude and direction of the net force on Q.



1
Expert's answer
2022-03-17T09:35:49-0400
"F_{net} = 2 | F | \cos \alpha"

"F=k\frac{qQ}{r^2}"


"\cos\alpha=\frac{4}{5}"


"F_{net}=2*9*10^9*\frac{4.0*10^{-6}*5.0*10^{-6}}{0.5^2}*\frac{4}{5}"

"F_{net}=\rm 1.15\: N"

Force is directed in the +x direction.